using the balanced equation
6 KI + 8 HNO3 = 6 KNO3 + 2 NO + 3 I2 + 4 H2O

When the reaction produces 8.0 mol of NO, how many molecules of I2 will be produced?
Answer #1

From the balanced equation we can say that

2 mole of NO produces 3 mole of I2 so

8.0 mole of NO will produce

= 8.0 mole of NO *(3 mole of I2 / 2 mole of NO)

= 12.0 mole of I2

1 mole of I2 = 6.023*10^23 molecules so

12.0 mole of I2 = 7.23*10^24 molecules

Therefore, the number of molecules of I2 =  7.23*10^24 molecules